mbuchanan:

Dan, Eric, and Frank are the leading scorers on their basketball team. Over the course of their high school careers, Dan scored \frac{6}{10}of the points that the three boys scored. If Eric scored \frac{7}{10}of the remaining points, and Frank scored 120 points, how many points did Dan score? 200 points 360 points 480 points 600 points

jhonyy9:

so than i understand it right Dan scored 6/10 of the points that the 3 boys scored Eric scored 7/10 of the remaining points Franc scored 120 points let the total points scored note it by a P using these above wrote we can write an equation \[\frac{ 6 }{ 10 }P +\frac{ 7 }{ 10 }\left( P -\frac{ 6 }{ 10 }P \right) +120 = P\] so now just you need solve this equation for P ,what will be the total points scored and after you get the value of P so just calcule the 6/10 value of P and in this way you ve get the right correct answer hope helped understandably easy